OpenStudy (anonymous):

Find the values of A and B such as (see reply). Algebraic Equivalence. Show work.

OpenStudy (anonymous):

\[\frac{ 1 }{ (x + 1)(x-2) } = \frac { A}{x+1} + \frac { B}{x-2}\] I tried to solve this, however, since Both sides are equalivant, so they seem to cancel each other out to a point A + B = 0. However this is not the correct answer.

OpenStudy (anonymous):

So from what I can gather, is that A = -B, But I havent got an idea how to find out the value.

OpenStudy (anonymous):